During the spin cycle of your clothes washer, the tub rotates at a steady angular velocity of 36.5 rad/s. Find the angular displacement of the tub during a spin of 68.1 s, expressed both in radians and in revolutions.
Answer #1

The equation of angular displacement of the tub is,

$$ \Delta \theta=\omega t $$

Here, \(\omega\) is angular velocity and \(t\) is time.

$$ \Delta \theta=(36.5 \mathrm{rad} / \mathrm{s})(68.1 \mathrm{~s})=2485.65 \mathrm{rad} $$

The angular displacement of the tub in radians is \(2485.65 \mathrm{rad}\)

Convert units of angular displacement from rad to rev.

$$ \begin{aligned} \Delta \theta &=2485.65 \mathrm{rad}\left(\frac{\mathrm{rev}}{2 \pi \mathrm{rad}}\right) \\ &=395.6 \mathrm{rev} \end{aligned} $$

The angular displacement of the tub in revolutions is \(395.6 \mathrm{rev}\)